# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ Abstract
AI-generated faces have enriched human life, such as entertainment, education, and art. However, they also pose misuse risks. Therefore, detecting AI-generated faces becomes crucial, yet current detectors show biased performance across different demographic groups. Mitigating biases can be done by designing algorithmic fairness methods, which usually require demographically annotated face datasets for model training. However, no existing dataset comprehensively encompasses both demographic attributes and diverse generative methods, which hinders the development of fair detectors for AI-generated faces. In this work, we introduce the AI-Face dataset, the first million-scale demographically annotated AI-generated face image dataset, including real faces, faces from deepfake videos, and faces generated by Generative Adversarial Networks and Diffusion Models. Based on this dataset, we conduct the first comprehensive fairness benchmark to assess various AI face detectors and provide valuable insights and findings to promote the future fair design of AI face detectors. Our AI-Face dataset and benchmark code are publicly available at https://github.com/Purdue-M2/AI-Face-FairnessBench.
Figure 1: Overview of AI-Face dataset. Each face has three demographic annotations with uncertainty scores.
@@ 1 Introduction
However, one major issue existing in current AI face detectors [14, 15, 16, 17] is biased detection (i.e., unfair detection performance among demographic groups [18, 19, 20, 21]). Mitigating biases can be done by designing algorithmic fairness methods, but they usually require demographically annotated face datasets for model training. For example, works like  [20, 21] have made efforts to enhance fairness in the detection based on A-FF++ [19] and A-DFD [19]. However, both datasets are limited to containing only faces from deepfake videos, which could cause the trained models not to be applicable for fairly detecting faces generated by GANs and DMs. Although a few datasets (e.g., GenData [22]) cover GAN and DM faces, their demographic annotations are not comprehensive. Most importantly, no existing dataset is diverse enough in generation methods to develop AI face detectors that can cope with rapidly evolved generative models. These limitations of existing datasets hamper the development of fair technologies for detecting AI-generated faces.
@@
Moreover, benchmarking fairness provides a direct method to uncover prevalent and unique fairness issues in recent AI-generated face detection. However, there is a lack of a comprehensive benchmark to estimate the fairness of existing AI face detectors. Existing benchmarks [23, 24, 25, 26] primarily assess utility, neglecting systematic fairness evaluation. One study [18] does evaluate fairness in detection models, but their examination is only based on deepfake video datasets using a few outdated detectors. Detectors’ fairness performance on GAN faces and DM faces has not been extensively explored. The absence of a comprehensive fairness benchmark impedes a thorough understanding of the fairness behaviors of recent AI face detectors and obscures the research path for detector fairness guarantees.
@@
In this work, we build the first million-scale demographically annotated AI-generated face image dataset: AI-Face (see Fig. 1). The face images are collected from various public datasets, including the real faces that are usually used to train AI face generators, faces from deepfake videos, and faces generated by GANs and DMs. Each face is demographically annotated with an uncertainty score on each predicted demographic attribute by our designed Contrastive Language-Image Pretraining (CLIP) [27]-based lightweight annotator. To improve the quality of annotations, we recruit three humans to correct annotations with high uncertainty scores manually. Next, we conduct the first comprehensive fairness benchmark on our dataset to estimate the fairness performance of 12 representative detectors coming from four model types. Our benchmark exposes common and unique fairness challenges in recent AI face detectors, providing essential insights that can guide and enhance the future design of fair AI face detectors. Our contributions are as follows:
@@
  • We build the first comprehensive million-scale demographically annotated AI-generated face Dataset by leveraging our developed lightweight annotator with human correction.
@@
  • Based on our experiments and observations, we summarize the unsolved questions and offer valuable insights within this research domain, setting the stage for future investigations.
@@ 2 Background and Motivation
AI-generated Faces and Biased Detection. AI-generated face images, created by advanced AI technologies, are visually difficult to discern from real ones, see Fig. 1. They can be summarized into three categories: 1) Deepfake Videos. Initiated in 2017 [13], these use face-swapping techniques with a variational autoencoder to replace a face in a target video with one from a source [3, 4]. Note that our paper focuses solely on images extracted from videos. 2) GAN-generated Faces. Post-2017, Generative Adversarial Networks (GANs) [28] like StyleGANs [6, 7, 8] have significantly improved generated face realism. 3) DM-generated Faces. Diffusion models (DMs), emerging in 2021, generate detailed faces from textual descriptions and offer greater controllability. Tools like Midjourney [29] and DALLE2 [30] facilitate customized face generation. While these AI-generated faces can enhance visual media and creativity [10], they also pose risks, such as being misused in social media profiles [31, 32]. Therefore, numerous studies focus on detecting AI-generated faces [14, 15, 16, 17], but current detectors often show performance disparities among demographic groups like race and gender [18, 19, 20, 21]. This bias can lead to unfair targeting or exclusion, undermining trust in detection models. Recent efforts [20, 21] aim to enhance fairness in deepfake detection but mainly address deepfake videos, overlooking biases in detecting GAN and DM-generated faces.
Dataset Face Images Generation Category #Generation Methods Source of Real Images Demographic Annotation
#Real #Fake Deepfake Videos GAN DM Gender Race Age
A-FF++ [19] 29.8K 149.1K 5 YouTube
A-DFD [19] 10.8K 89.6K 5 Self-Recording
A-DFDC [19] 54.5K 52.6K 8 Self-Recording
A-Celeb-DF-v2 [19] 26.3K 166.5K 1 Self-Recording
A-DF-1.0 [19] 870.3K 321.5K 1 Self-Recording
DF-1.0 [33] 2.9M 14.7M 1 Self-Recording
DeePhy [34] 1K 50.4K 3 YouTube
DF-Platter [35] 392.3K 653.4K 3 YouTube
GenData [22] - 20K 3 CelebA [36]
Ours 866K 1.2M 37
FFHQ [6], CASIA-WebFace [37], CelebA [36]
IMDB-WIKI [38], real from FF++ [2],
DFDC [39], DFD [40],Celeb-DF-v2 [41]
Table 1: Quantitative comparison of existing AI-generated face datasets and ours.
@@
The Related Existing Datasets. Current AI-generated facial datasets with demographic annotations are limited in size, generation categories, methods, and annotations, as illustrated in Table 1. For instance, A-FF++, A-DFD, A-DFDC, and A-Celeb-DF-v2 [19] are deepfake video datasets with fewer than one million images. Datasets like DF-1.0 [33] and DF-Platter [35] lack comprehensive demographic annotations. Additionally, existing datasets offer limited generation methods. These limitations hinder the development of fairer AI face detectors, motivating us to build a million-scale demographically annotated AI-Face dataset.
Existing Benchmarks Category Scope of Benchmark
Deepfake Videos GAN DM Utility Fairness
DeepfakeBench [25]
Lin et al. [24]
Le et al. [26]
CDDB [23]
Loc et al. [18]
Ours
Table 2: Comparison with existing AI-generated face detection benchmarks.
@@
Benchmark for Detecting AI-generated Faces. Benchmarks are essential for evaluating AI-generated face detectors under standardized conditions. Existing benchmarks, as shown in Table 2, mainly focus on detectors’ utility, often overlooking fairness [23, 24, 25, 26]. Only Loc et al. [18] examined detector fairness. However, their study focused only on deepfake video datasets, not on GAN- and DM-generated faces. This motivates us to conduct a comprehensive benchmark to evaluate AI face detectors’ fairness.
3 The Demographically Annotated AI-Face Dataset
To address the prohibitive time consuming of manual annotation, we introduce two phases to build our dataset: Annotator Development and Demographically Annotation Generation, as shown in Fig. 2.
3.1 Phase 1: Annotator Development
Problem Definition. There are existing online software (e.g., Face++ [42]) and open-source tools (e.g., InsightFace [43]) for face attribute prediction. However, they fall short of our task due to two reasons: 1) They are mostly designed for face recognition and trained on datasets of real face images but lack generalization capability for annotating AI-generated face images. 2) They do not provide uncertainty scores for their predictions that can be used to identify mispredicted samples for further annotation correction. Given a training dataset $\mathbb{D}=\{(X_{i},G_{i},A_{i},R_{i})\}_{i=1}^{n}$ with size $n$, where $X_{i}$, $G_{i}\in\{Female,Male\}$, $A_{i}\in\{Young,Middle\text{-}aged,Senior,Others\}$, and $R_{i}\in\{Asian,White,Black,Others\}$ represent the $i$-th face image, and its gender, age, and race labels/attributes, respectively. Our goal is to design a lightweight, generalizable annotator based on $\mathbb{D}$ to predict facial demographic attributes with uncertainty scores for each face image in our dataset.
@@ 3.1 Phase 1: Annotator Development
Figure 2: Generation pipeline of our Demographically Annotated AI-Face Dataset. First, we develop a lightweight facial attribute annotator trained on the VGGFace2 dataset. Second, we collect and filter face images from Deepfake Videos, GAN-generated faces, and DM-generated faces found in public datasets. Our AI-Face annotator then predicts facial attributes with uncertainty scores for each image. Third, samples with high uncertainty are manually reviewed and corrected by three humans to improve annotation quality.
Annotator. Architecture: We utilize CLIP [27] for its strong zero-shot and few-shot learning capabilities. Leveraging CLIP’s pre-training on diverse datasets, we create a lightweight annotator for facial images. Our annotator employs a frozen pre-trained CLIP ViT L/14 [44] as a feature extractor $\mathbf{E}$ followed by a trainable 3-layer Multilayer Perceptron (MLP) as a multi-task (i.e., gender, age, and race prediction) classifier parameterized by $\theta$. Loss: For each image $X_{i}$, its feature $f_{i}$ is obtained through $f_{i}=\mathbf{E}(X_{i})$ and then is fed into the MLP multi-task classifier with conventional classification losses for face attribute prediction. The learning objective is formulated as: $\mathcal{L}(\theta)=C(\widetilde{h}(f_{i}),G_{i})+C(\overline{h}(f_{i}),A_{i})+C(\widehat{h}(f_{i}),R_{i})$, where $C(\cdot,\cdot)$ represents the (binary) cross-entropy (CE) loss. $\widetilde{h}$, $\overline{h}$, and $\widehat{h}$ represent the classification heads for gender, age, and race, respectively. Optimization: Traditional optimization methods like stochastic gradient descent can lead to poor model generalization due to sharp loss landscapes with multiple local and global minima. To address this, we use Sharpness-Aware Minimization (SAM) [45] to enhance our annotator’s generalization by flattening the loss landscape. Specifically, flattening is attained by determining the optimal $\epsilon^{*}$ for perturbing model parameters $\theta$ to maximize the loss, formulated as: $\epsilon^{*}=\arg\max_{\|\epsilon\|_{2}\leq\gamma}{\mathcal{L}}\textbf{(}\theta+\epsilon\textbf{)}\approx\arg\max_{\|\epsilon\|_{2}\leq\gamma}\epsilon^{\top}\nabla_{\theta}\mathcal{L}=\gamma\texttt{sign}(\nabla_{\theta}\mathcal{L})$, where $\gamma$ controls the perturbation magnitude. This is approximated using a first-order Taylor expansion, assuming $\epsilon$ is small. The final equation is obtained by solving a dual norm problem, where sign represents a sign function and $\nabla_{\theta}\mathcal{L}$ being the gradient of $\mathcal{L}$ with respect to $\theta$. As a result, the model parameters are updated by solving: $\min_{\theta}\mathcal{L}\textbf{(}\theta+\epsilon^{*}\textbf{)}$.

### v3

@@ Abstract
AI-generated faces have enriched human life, such as entertainment, education, and art. However, they also pose misuse risks. Therefore, detecting AI-generated faces becomes crucial, yet current detectors show biased performance across different demographic groups. Mitigating biases can be done by designing algorithmic fairness methods, which usually require demographically annotated face datasets for model training. However, no existing dataset encompasses both demographic attributes and diverse generative methods simultaneously, which hinders the development of fair detectors for AI-generated faces. In this work, we introduce the AI-Face dataset, the first million-scale demographically annotated AI-generated face image dataset, including real faces, faces from deepfake videos, and faces generated by Generative Adversarial Networks and Diffusion Models. Based on this dataset, we conduct the first comprehensive fairness benchmark to assess various AI face detectors and provide valuable insights and findings to promote the future fair design of AI face detectors. Our AI-Face dataset and benchmark code are publicly available at https://github.com/Purdue-M2/AI-Face-FairnessBench.
footnotetext: Corresponding author
@@ 1 Introduction
Figure 1: Comparison between AI-Face and other datasets in terms of demographic annotation, generation category, and the number of generation methods. ‘DF’, ‘GAN’, and ‘DM’ stand for Deepfake Videos, Generative Adversarial Networks, and Diffusion Models.
  Face Images Generation Category Demographic Annotation
Dataset Year #Real #Fake Deepfake Videos GAN DM #Generation Methods Source of Real Images Skin Tone Gender Age
DF-1.0 [14] 2020 2.9M 14.7M 1 Self-Recording
DeePhy [15] 2022 1K 50.4K 3 YouTube
DF-Platter [16] 2023 392.3K 653.4K 3 YouTube
GenData [17] 2023 - 20K 3 CelebA [18]
A-FF++ [19] 2024 29.8K 149.1K 5 YouTube
A-DFD [19] 2024 10.8K 89.6K 5 Self-Recording
A-DFDC [19] 2024 54.5K 52.6K 8 Self-Recording
A-Celeb-DF-v2 [19] 2024 26.3K 166.5K 1 Self-Recording
A-DF-1.0 [19] 2024 870.3K 321.5K 1 Self-Recording
AI-Face 2025 400K 1.2M 37
FFHQ [6], IMDB-WIKI [20],
real from FF++ [2], DFDC [21],
DFD [22],Celeb-DF-v2 [23]
Table 1: Quantitative comparison of existing datasets with ours on demographically annotated AI-generated faces.
@@
However, one major issue existing in current AI face detectors [24, 25, 26, 27] is biased detection (i.e., unfair detection performance among demographic groups [28, 19, 29, 30]). Mitigating biases can be done by designing algorithmic fairness methods, but they usually require demographically annotated face datasets for model training. For example, works like  [29, 30] have made efforts to enhance fairness in the detection based on A-FF++ [19] and A-DFD [19]. However, both datasets are limited to containing only faces from deepfake videos, which could cause the trained models not to be applicable for fairly detecting faces generated by GANs and DMs. While some datasets (e.g., GenData [17], DF40 [31]) include GAN and DM faces, they either lack demographic annotations or provide only limited demographic attributes. Most importantly, no existing dataset offers sufficient diversity in generation methods while also providing demographic labels. A comparison of existing datasets is shown in Fig. 1. These limitations of existing datasets hamper the development of fair technologies for detecting AI-generated faces.
@@
Moreover, benchmarking fairness provides a direct method to uncover prevalent and unique fairness issues in recent AI-generated face detection. However, there is a lack of a comprehensive benchmark to estimate the fairness of existing AI face detectors. Existing benchmarks [32, 33, 34, 35] primarily assess utility, neglecting systematic fairness evaluation. Two studies [28, 36] do evaluate fairness in detection models, but their examination is based on a few outdated detectors. Furthermore, detectors’ fairness reliability (e.g., robustness with test set post-processing, fairness generalization) has not been assessed. The absence of a comprehensive fairness benchmark impedes a thorough understanding of the fairness behaviors of recent AI face detectors and obscures the research path for detector fairness guarantees.
@@
In this work, we build the first million-scale demographically annotated AI-generated face image dataset: AI-Face. The face images are collected from various public datasets, including the real faces that are usually used to train AI face generators, faces from deepfake videos, and faces generated by GANs and DMs. Each face is demographically annotated by our designed measurement method and Contrastive Language-Image Pretraining (CLIP) [37]-based lightweight annotator. Next, we conduct the first comprehensive fairness benchmark on our dataset to estimate the fairness performance of 12 representative detectors coming from four model types. Our benchmark exposes common and unique fairness challenges in recent AI face detectors, providing essential insights that can guide and enhance the future design of fair AI face detectors. Our contributions are as follows:
@@
  • We build the first million-scale demographically annotated AI-generated face dataset by leveraging our designed measure and developed lightweight annotator.
@@
  • Based on our experiments, we summarize the unsolved questions and offer valuable insights within this research field, setting the stage for future investigations.
@@ 2 Background and Motivation
AI-generated Faces and Biased Detection. AI-generated face images, created by advanced AI technologies, are visually difficult to discern from real ones. They can be summarized into three categories: 1) Deepfake Videos. Initiated in 2017 [13], these use face-swapping and face-reenactment techniques with a variational autoencoder to replace a face in a target video with one from a source [3, 4]. Note that our paper focuses solely on images extracted from videos. 2) GAN-generated Faces. Post-2017, Generative Adversarial Networks (GANs) [38] like StyleGANs [6, 7, 8] have significantly improved generated face realism. 3) DM-generated Faces. Diffusion models (DMs), emerging in 2021, generate detailed faces from textual descriptions and offer greater controllability. Tools like Midjourney [39] and DALLE2 [40] facilitate customized face generation. While these AI-generated faces can enhance visual media and creativity [10], they also pose risks, such as being misused in social media profiles [41, 42]. Therefore, numerous studies focus on detecting AI-generated faces [24, 25, 26, 27], but current detectors often show performance disparities among demographic groups [28, 19, 29, 30]. This bias can lead to unfair targeting or exclusion, undermining trust in detection models. Recent efforts [29, 30] aim to enhance fairness in deepfake detection but mainly address deepfake videos, overlooking biases in detecting GAN- and DM-generated faces.
@@
The Existing Datasets. Current AI-generated facial datasets with demographic annotations are limited in size, generation categories, methods, and annotations, as illustrated in Table 1. For instance, A-FF++, A-DFD, A-DFDC, and A-Celeb-DF-v2 [19] are deepfake video datasets with fewer than one million images. Datasets like DF-1.0 [14] and DF-Platter [16] lack various demographic annotations. Additionally, existing datasets offer limited generation methods. These limitations hinder the development of fair AI face detectors, motivating us to build a million-scale demographically annotated AI-Face dataset.
@@
Benchmark for Detecting AI-generated Faces. Benchmarks are essential for evaluating AI-generated face detectors under standardized conditions. Existing benchmarks, as shown in Table 2, mainly focus on detectors’ utility, often overlooking fairness [32, 33, 34, 35, 31]. Loc et al. [28] and CCv1 [36] examined detector fairness. However, their study did not have an analysis on DM-generated faces and only measured bias between groups in basic scenarios without considering fairness reliability under real-world variations and transformations. This motivates us to conduct a comprehensive benchmark to evaluate AI face detectors’ fairness.
The Definition of Demographic Categories. Demography-related labels are highly salient to measuring bias. Following prior works [43, 44, 45, 36, 46, 47], we will focus on three key demographic categories: Skin Tone, Gender, and Age, in this work. For skin tone, this vital attribute spans a range from pale to dark. We use the Monk Skin Tone scale [48], specifically designed for computer vision applications. For gender, we adopt binary categories (i.e., Male and Female), following practices by many governments [49, 50] and facial recognition research [45, 51, 52], based on sex at birth. For age, using definitions from the United Nations [53] and Statistics Canada [54], we define five age groups: Child (0-14), Youth (15-24), Adult (25-44), Middle-age Adult (45-64), and Senior (65+). More discussion is in Appendix A.
3 AI-Face Dataset
This section outlines the process of building our demographically annotated AI-Face dataset (see Fig. 2), along with its statistics and annotation quality assessment.
  Category Scope of Benchmark
Fairness
Existing Benchmarks Year Deepfake Videos GAN DM Utility General Reliability
Loc et al. [28] 2021
CCv1 [36] 2021
DeepfakeBench [34] 2023
CDDB [32] 2023
Lin et al. [33] 2024
Le et al. [35] 2024
DF40 [31] 2024
Ours 2025
Table 2: Comparison of existing AI-generated face detection benchmarks and ours. Fairness ‘General’ means fairness evaluation under default/basic settings. Fairness ‘Reliability’ measures fairness consistency across dynamic scenarios (e.g., post-processing).
@@ 3 AI-Face Dataset
Figure 2: Generation pipeline of our Demographically Annotated AI-Face Dataset. First, we collect and filter face images from Deepfake Videos, GAN-generated faces, and DM-generated faces found in public datasets. Second, we perform skin tone, gender, and age annotation generation. Skin tone is estimated by combining facial landmark detection with color analysis to generate the corresponding annotation. For gender and age, we develop annotators trained on the IMDB-WIKI dataset [20], then use them to predict attributes for each image.
3.1 Data Collection
We build our AI-Face dataset by collecting and integrating public real and AI-generated face images sourced from academic publications, GitHub repositories, and commercial tools. We strictly adhere to the license agreements of all datasets to ensure that they allow inclusion in our datasets and secondary use for training and testing. More details are in Appendix B.1. Specifically, the fake face images in our dataset originate from 4 Deepfake Video datasets (i.e., FF++ [2], DFDC [21], DFD [22], and Celeb-DF-v2 [23]), generated by 10 GAN models (i.e., AttGAN [55], MMDGAN [56], StarGAN [55], StyleGANs [55, 57, 58], MSGGAN [56], ProGAN [59], STGAN [56], and VQGAN [60]), and 8 DM models (i.e., DALLE2 [61], IF [61], Midjourney [61], DCFace [62], Latent Diffusion [63], Palette [64], Stable Diffusion v1.5 [65], Stable Diffusion Inpainting [65]). This constitutes a total of 1,245,660 fake face images in our dataset. We include 6 real source datasets (i.e., FFHQ [6], IMDB-WIKI [20], and real images from FF++ [2], DFDC [21], DFD [22], and Celeb-DF-v2 [23]). All of them are usually used as a training set for generative models to generate fake face images. This constitutes a total of 400,885 real face images in our dataset. In general, our dataset contains 28 subsets and 37 generation methods (i.e., 5 in FF++, 5 in DFD, 8 in DFDC, 1 in Celeb-DF-v2, 10 GANs, and 8 DMs). For all images, we use RetinaFace [66] for detecting and cropping faces.
3.2 Annotation Generation
3.2.1 Skin Tone Annotation Generation
Skin tone is typically measured using an intuitive approach [67, 68], without requiring a predictive model. Inspired by [67], we developed a method to estimate skin tone using the Monk Skin Tone (MST) Scale [48] (including 10-shade scales: Tone 1 to 10) by combining facial landmark detection with color analysis. Specifically, utilizing Mediapipe’s FaceMesh [69] for precise facial landmark localization, we isolate skin regions while excluding non-skin areas such as the eyes and mouth. Based on the detected landmarks, we generate a mask to extract skin pixels from the facial area. These pixels are then subjected to K-Means clustering [70] (we use K$=3$ in practice) to identify the dominant skin color within the region of interest. The top-1 largest color cluster is mapped to the closest tone in the MST Scale by calculating the Euclidean distance between the cluster centroid and the MST reference colors in RGB space.
3.2.2 Gender and Age Annotation Generation
For generating gender and age annotations, the existing online software (e.g., Face++ [71]) and open-source tools (e.g., InsightFace [72]) can be used for the prediction. However, they fall short in our task due to two reasons: 1) They are mostly designed for face recognition and trained on datasets of real face images but lack generalization capability for annotating AI-generated face images. 2) Their use may introduce bias into our dataset, as they are typically designed and trained without careful consideration of bias and imbalance in the training set. See Appendix  B.3 for our experimental study on these tools. To this end, we have to develop our specific annotators to predict gender and age annotations for each image in our dataset.